### v1

@@ Abstract
Quantum systems have started to emerge as a disruptive technology and enabling platforms – exploiting the principles of quantum mechanics via programmable quantum bits (QuBits) – to achieve quantum supremacy in computing. Academic research, industrial projects (e.g., Amazon Braket, IBM Qiskit), and consortiums like ‘Quantum Flagship’ are striving to develop practically capable and commercially viable quantum computing (QC) systems and technologies. Quantum Computing as a Service (QCaaS) is viewed as a solution attuned to the philosophy of service-orientation that can offer QC resources and platforms, as utility computing, to individuals and organisations who do not own quantum computers. This research investigates a process-centric and architecture-driven approach to offer a software engineering perspective on enabling QCaaS - a.k.a quantum service-orientation. To conduct this research, we employed a two-phase research method comprising (a) a systematic mapping study and (b) an architecture-based development, first to identify the phases of the quantum service development life cycle and subsequently to integrate these phases into a reference architecture that supports QCaaS. The systematic mapping study (SMS) process retrieved a collection of potentially relevant research literature and based on a multi-step selection and qualitative assessment, we selected 41 peer-reviewed studies as published evidence to answer three research questions (RQs). The RQs investigate (i) demographic details in terms of frequency, types, and trends of research, (ii) phases of quantum service development lifecycle to derive a reference architecture for conception, modeling, assembly, and deployment of services, and (iii) emerging trends that can shape future research on software engineering for QCaaS. The results identify a 4-phased development lifecycle along with quantum significant requirements (QSRs), various modeling notations, catalogue of patterns, programming languages, and deployment platforms that can be integrated in a layered reference architecture to engineer QCaaS. Emerging and future research efforts focus on topics such as classifying the QSRs, enabling model-driven or low-code quantum service engineering, and integrating large language models in the quantum service development process. In the broader perspective of quantum software engineering (QSE), the results of this SMS specifically focus on quantum service-orientation by providing empirically grounded findings, highlighting patterns for knowledge reuse, and offering a reference architecture as a blueprint to engineer software for QCaaS.
@@ 1 Introduction
Quantum Computers (QCs) aim to disrupt the status-quo of computing – replacing traditional systems and platforms that are driven by digital circuits and modular applications – with hardware and software that operate on the principle of quantum mechanics [1]. QCs that rely on quantum mechanics can exploit quantum circuits via programmable quantum bits (QuBits) to control quantum gates (QuGates) for achieving ‘quantum computational supremacy’ over their classical counterparts, i.e., digital computing systems [2]. QC systems are in a phase of continuous evolution, and despite being in a state of their infancy due to hardware limitations or lack of software ecosystem(s), such systems have started to computationally outperform classical computers in certain application areas such as quantum information processing, bio-inspired computing, and simulation of quantum mechanics [2]. Academic research [1] [3] and industrial initiatives led by technology giants such as IBM, Google, and Microsoft [4] [5] are striving to achieve strategic advantages associated with quantum systems and software technologies in a so-called ‘race to quantum economy’. A recent report presented at the World Economic Forum titled State of Quantum Computing: Building a Quantum Economy highlights that by the year 2022, public and private investments in quantum computing technologies totalled \$35.5 billion [6] and global investments in QCs reached up to \$55 billion in 2024 [7]. Despite the strategic capabilities that quantum supremacy in computing can offer, programming, operationalising, and maintaining a quantum computer represent a complex and radically distinct engineering paradigm compared to the challenges of classical computing and software engineering [1]. QC specific challenges may include quantum domain engineering, circuit modelling, algorithm design, and classic-quantum hybrid system deployment. These challenges arise from a lack of systematic processes, limited tool support, and a scarcity of human expertise in quantum software development. To address these challenges and support QC innovations, state-funded projects and global consortiums are proactively funding initiatives such as the Quantum Flagship [8] and National Quantum Initiative [9] to develop software ecosystems, networking technologies, and human expertise for the alleged quantum leap in computing [10].
@@
Research context and motivation: Service-oriented systems support the utility computing model, enabling individuals and organisations to access multiple computing services on a pay-per-use basis without owning or maintaining them [11]. Since their early adoption, pay-per-use service-driven systems have now grown from data storage, video streaming, and entertainment, to resource-sharing applications that represent a multi-billion dollar industry in service economies [12][13]. The ‘as-a-Service’ (aaS) model has provided the impetus to wide-scale adoption of service systems that can offer a plethora of computing services including but not limited to storage, computation, infrastructure, platform, and software to end-users [12]. The aaS model can enable users and developers who can exploit the QC platforms (e.g., processors, memory, simulators) offered by quantum vendors, such as Amazon, Google, and IBM [14]. Quantum Computing as a Service (QCaaS) as shown in Figure 1 is a recent and quantum-specific genre of aaS model that is built on the philosophy of service-orientation of QC, i.e., pay-per-shot at QC resources instead of owning, programming, and/or maintaining quantum computers [15]. Quantum vendors view pay-per-shot as an opportunistic business model to generate revenue streams from their QC infrastructures, where a shot is a single execution of a quantum task/algorithm on a quantum processing unit (QPU). In simple terms, we define QCaaS as a ‘quantum-specific variation of the aaS model that offers quantum hardware and software resources such as processor, memory, algorithms, and simulations that can be utilised by users (via client nodes) by connecting to a remotely deployed quantum computer (i.e., server)’. The QCaaS model can alleviate the need to own or maintain quantum computers and can help software and service developers to rely on existing knowledge and best practices to develop software services that can be executed on QC platforms [16]. Quantum software services allow developers to encapsulate data and computations in loosely coupled, fine-grained modules to execute tasks on QC platforms [15]. As shown in Figure 1, service developers can leverage reusable knowledge and best practices from classical service engineering—such as microservices, architecture, and patterns—to design quantum services efficiently and effectively. The developed quantum services can enable quantum users, i.e., service requesters, to utilise QC resources offered by quantum vendors, i.e., service providers via pay-per-shot QCaaS model. Without a systematic and structured approach and a lack of expertise in QC software quantum service development risks include inefficient Qubit utilisation, suboptimal quantum-classical task allocation, error-prone circuit design etc., ultimately constraining the success and sustainability of QCaaS [17].
@@
Software Engineering for QCaaS: Quantum software engineering, as a quantum-specific genre of classical software engineering, leverages established concepts such as architecture models as system blueprints and patterns as reusable knowledge to support software service developers in engineering solutions for QCaaS [18]. By applying SE best practices, developers can, for instance, utilise the quantum-classical split pattern to design hybrid applications, map architectural components to quantum circuits, and enable model-driven quantum code generation [19] [20]. In this research, we have used an SE-focused two-phase research method that relies on a systematic mapping study to investigate existing literature and follow the architecture-based development for an incremental design, implementation, and deployment of quantum software services. Systematic mapping studies (SMS) use an evidence-based software engineering approach to systematically identify, analyse, synthesise, and document research trends on a given topic [21]. The academic community aims to exploit QC platforms to advance state-of-the-art and develop prototypes to support research [14], while the vendors pursue a revenue stream as well as validation and testing of their under-development quantum platforms, offered as a service [15]. Synergising research and development on quantum computing with service-orientation can allow researchers and developers to leverage existing design principles, patterns, architectural styles, and modelling languages of service computing to offer QC as a service [16]. In the context of QSE, this research synthesises the available evidence reflected via published research to provide a software engineering perspective on architecting, implementing, and operationalising QCaaS. We conducted this study to answer three research questions that investigate (i) demographics: trends, types, frequency and progression of research (ii) quantum service lifecycle: that can be integrated in a layered reference architecture for conceptualising, modeling, assembling, and deploying services, and (ii) emerging research: to highlight topics for potentially futuristic research on QCaaS. The results indicate that in quantum service-orientation, QSRs, both functional aspects and quality attributes, guide the four-phase service development lifecycle. Software modeling languages (e.g., UML, circuit diagrams) and patterns (e.g., API gateway, quantum-classic split) can be applied within the proposed layered architecture of QCaaS via quantum-classical hybrid computing. Emerging trends include QSR classification, model-driven development, empirically derived patterns, and defined human roles in QCaaS. The main contributions and implications of this research are to:
@@
This research represents one of the pioneering efforts to rely on evidence-based software engineering guidelines to investigate QCaaS and its implications on QSE. The study results can have implications for academic researchers who can rely on mapping of existing research and vision for future work to research and develop QCaaS [1][8]. The results can also guide practitioners who can rely on academic references about patterns (reusability), notations (representation), and implementation (prototyping) to architect QCaaS [14].
Technical context of QC and QSE is presented in Section 2. Research method is discussed in Section 3. Results of the study to answer the RQs are presented from Section 4 - 7. Review of the existing research is presented in Section 8. Implications and validity threats of the study are discussed in Section 9. Conclusions and future work are presented in Section 10.
@@ 2 Research Context
This section contextualises the core concepts used in our study and elaborates them based on illustrations in Figure 2. First, we overview the fundamental elements of quantum computing and the role of quantum software engineering to operationalise QC systems in Section 2.1, as shown in Figure 2 a). Second, we explain service-orientation from quantum software engineering perspective in Section 2.2 via Figure 2 b). The concepts and terminologies introduced in this section are used throughout the paper.
Figure 2: An Overview of the Quantum Service-Orientation
@@ 2.1.1 Quantum Computing Systems
We discuss the quantum computing system from both the hardware and software perspective as shown in Figure 2 a). Figure 2 is adopted from [19] and extended to highlight software architecting in the context of quantum service-orientation. We briefly overview a QC system that comprises of quantum hardware and software elements as shown in Figure 2 a). Fundamental to enabling quantum computation are Quantum Bits (QuBits) that represent the basic unit of quantum information processing by manipulating Quantum Gates (QuGates) [2] [10]. Traditional Binary Digits (Bits) in classical systems (i.e., digital computers) are represented as [1, 0] where 1 represents the computation state as On and 0 represents the state as Off to manipulate binary gates in digital circuits. In comparison, a QuBit represents a two-state quantum computer expressed as $|0\rangle$ and $|1\rangle$. The state of a single QuBit can be expressed as $|0\rangle=\begin{bmatrix}1\\ 0\end{bmatrix}$ and $|1\rangle=\begin{bmatrix}0\\ 1\end{bmatrix}$ and quantum superposition allows a QuBit to attain a liner combination of both states:
@@
According to Figure 2 a), we distinguish between a Bit and QuBit such that a Bit can take a value as either ‘Off:0’ or ‘On:1’ with 100% probability. In comparison, a QuBit can be in a state of $|0\rangle$ or $|1\rangle$ or in a superposition state with 50% $|0\rangle$ and 50% $|1\rangle$. In addition, two QuBits can be entangled such that entangled QuBits are linked in a way that observing (i.e., measuring) one of the QuBits, can reveal the state of the other QuBit. Extended details about QuBits and QuGates to develop and operate the QC systems are reported in studies like [1] and [14]. To utilise the quantum computing resources, such as quantum processor and memory, there is a need for control software that can program QuBits to manage QuGates of a QC system. To support systems and services for quantum computing, there is a need for systematic processes that incorporates tools, models, patterns, frameworks, and human expertise as engineering practices to develop quantum software [22] [23].
@@ 2.1.2 Quantum Software Engineering
QSE has emerged as a quantum specific genre of classical software engineering and it focuses on unifying the principles and practices of quantum mechanics and practices of software engineering - to design, develop, validate, and evolve quantum age software systems, services, and applications [24] [25]. QSE empowers the role of software designers and developers who can exploit processes that support structured development, architectural models for design visualisation, and patterns as best practices to engineer software that can be executed on QCs. Quantum software systems rely on quantum source code compilers that allow quantum algorithm designers and programmers to develop and execute software for quantum computers. For example, a programmer can use a quantum programming language, such as Q# (by Microsoft) or Qiskit (by Google) and use a quantum compiler to enable programable quantum computations [4] [25]. Software systems capable of managing and controlling quantum hardware find applications in areas including, but not limited to, quantum cryptography, bio-inspired computing, and quantum information processing [1]. However, the scarcity of quantum hardware, the shortage of quantum software professionals, and the high costs associated with owning or maintaining quantum computers are critical factors impeding the commercial viability of quantum computing [2] [3]. Vendors who offer QCaaS platforms view quantum service-orientation as an opportunistic business model that offers QC resources to customers as utility computing model [11] [14].
@@ 2.2.1 Services Computing
It follows the service-oriented architecture (SOA) style that allows service users to discover and utilise a multitude of available software services that encapsulate computing resources and applications offered by service vendors/providers [12]. Fundamental to the concept of service computing is software service that represents a self-contained, reusable, and platform-independent software module that encapsulates specific functionality or business logic, accessible over a network through standardised protocols, typically as part of SOA [26]. More specific types of services can be web services or microservices that enable Software as a Service (SaaS) model in which software applications are hosted and managed by a provider, accessed by users over the network via a web client, and offered on a subscription basis, eliminating the need for local installation and maintenance [15]. Figure 2 b) shows SOA style-based quantum servicing where a QC user (i.e., service requester) can utilise the QC resources offered by quantum vendors (i.e., service provider) by means of quantum services.

### v2

@@ Abstract
Quantum systems are emerging as a transformative technology, harnessing the principles of quantum mechanics through programmable quantum bits (QuBits) to deliver, in certain cases, superior computational capabilities termed as quantum supremacy in computing. In recent years, academic research, industry led projects (e.g., Amazon Braket, IBM Quantum) as well as international consortiums such as the ‘Quantum Flagship’ are aiming to develop solutions for commercially viable quantum computing (QC). In this context, Quantum Computing as a Service (QCaaS) represents a service-oriented solution that provides access to quantum computing resources and platforms as utility-based services, enabling individuals and organisations to use QC capabilities without owning or managing quantum hardware and software. This research employs a process-centric and architecture-driven approach to offer a software engineering perspective on enabling QCaaS – a.k.a quantum service-orientation. To conduct this research, we followed a two-phase research method comprising (i) a systematic mapping study and (ii) an architecture-based development, first to identify the phases of the quantum service development lifecycle and subsequently to integrate these phases into a reference architecture that supports QCaaS. The systematic mapping study (SMS) process retrieved a collection of potentially relevant research literature and based on a multi-step selection and qualitative assessment, we selected 41 peer-reviewed studies as published evidence to answer three research questions (RQs). The RQs investigate (a) demographic details in terms of frequency, types, and trends of research, (b) phases of quantum service development lifecycle to derive a reference architecture for conception, modeling, assembly, and deployment of services, and (c) emerging research trends that can shape future practices on software engineering for QCaaS. The results identify a 4-phased development lifecycle along with quantum significant requirements (QSRs), various modeling notations, catalogue of patterns, programming languages, and deployment platforms that can be integrated in a layered reference architecture to engineer QCaaS. Emerging and future research efforts focus on topics such as classifying the QSRs, enabling model-driven or low-code quantum service engineering, and integrating large language models in the quantum service development process. In the broader perspective of quantum software engineering (QSE), the results of this SMS specifically focus on quantum service-orientation by providing empirically grounded findings, highlighting patterns for knowledge reuse, and offering a reference architecture as a blueprint to engineer QCaaS.
@@ 1 Introduction
Quantum computing (QC) aims to revolutionise the existing computing paradigm by operating on the principles of quantum mechanics, unlike conventional systems that rely on digital circuits and modular applications [1]. QCs utilise quantum bits (QuBits), which can be programmed to manipulate quantum gates (QuGates) to perform computational tasks that, in certain cases, surpass the capabilities of classical digital computers – a phenomenon known as quantum supremacy in computing [2]. QC systems are continuously evolving, and although still in their early stages due to hardware limitations or lack of software ecosystem(s), such systems have started to computationally outperform classical computers in certain application areas, including quantum information processing, bio-inspired computing, and quantum simulations [2]. Academic research [1][3] and industrial efforts led by leading QC vendors such as IBM, Google, and Microsoft [4] [5] are competing to secure strategic advantages in quantum technologies, often described as the race to the quantum economy. According to the World Economic Forum’s report titled State of Quantum Computing: Building a Quantum Economy, public and private investments in quantum computing reached \$35.5 billion by 2022 [6], rising to \$55 billion approx. globally in 2024 [7]. Despite the strategic potential of quantum supremacy, programming, operating, and maintaining quantum computers presents a complex and fundamentally different engineering challenge from classical computing and software engineering [1]. QC specific challenges include but are not limited to quantum domain engineering, circuit modeling, algorithm design, and quantum-classic hybrid system deployment. These challenges stem from a lack of systematic processes, insufficient tool support, and a shortage of expertise in quantum software and system development. To address such challenges, state-funded projects and global initiatives such as the Quantum Flagship [8] and National Quantum Initiative [9] are investing in developing hardware components, software ecosystems, networking technologies, and skilled human resources for the anticipated quantum leap in computing [10].
@@
Research context and motivation: Service-oriented systems support the utility computing model, enabling individuals and organisations to access multiple computing services on a pay-per-use basis without owning or maintaining them [11]. Since their inception and early adoption, pay-per-use, service-driven systems have grown from data storage, video streaming, and entertainment to resource-sharing applications, forming a multi-billion-dollar segment of the service economy [12][13]. The ‘as-a-Service (aaS)’ model has driven the widespread adoption of platforms that provide end users with various computing services – including storage, computation, infrastructure, platforms, and software – all delivered through on-demand access [12]. The aaS model also enables developers and users to utilise QC platforms (processors, memory, simulators) provided by vendors such as Amazon, Google, and IBM [14]. Quantum Computing as a Service (QCaaS), conceptualised in Figure 1, represents a quantum-specific extension of the aaS model. It follows the service-oriented philosophy of pay-per-use, allowing access to quantum resources without the need to own, program, or maintain quantum computers [15]. For QC vendors, pay-per-shot, where a shot is a single execution of a quantum task on a quantum processing unit (QPU), serves as a revenue-generating business model. In simple terms, we define QCaaS as a ‘quantum-specific variation of the aaS model that offers quantum hardware and software resources such as processor, memory, algorithms, and simulators etc. that can be utilised by users (via client nodes) by connecting to a remotely deployed quantum computer (i.e., server)’. The QCaaS model can alleviate the need to own or maintain QCs, allowing software and service developers to build applications on QC platforms by leveraging existing knowledge and best practices [16]. Quantum software services allow developers to encapsulate data and computations in loosely coupled, fine-grained modules to execute tasks on QC platforms [15]. As shown in Figure 1, service developers can leverage reusable knowledge and best practices from classical service engineering such as microservices frameworks and service design patterns to implement quantum services efficiently and effectively. The developed quantum services can enable quantum users, i.e., service requesters, to utilise QC resources offered by quantum vendors, i.e., service providers via pay-per-shot QCaaS model.Without a systematic and structured approach, and in the absence of sufficient expertise in QC service development, there is are risks of inefficient QuBit utilisation, suboptimal quantum–classical task allocation, and error-prone circuit design, all of which can ultimately hinder the success and sustainability of QCaaS. [17].
@@
Software Engineering for QCaaS: Quantum software engineering, as a quantum-specific genre of classical software engineering, leverages established concepts such as architecture models as system blueprints and patterns as reusable knowledge to support software service developers in engineering solutions for QCaaS [18]. By applying SE best practices, developers can, for instance, utilise the quantum-classic split pattern to design hybrid applications, map architectural components to quantum circuits, and enable model-driven quantum code generation [19] [20]. In this research, we have used an SE-focused two-phase research method that relies on a systematic mapping study [21] to investigate existing literature and follow the architecture-based development [22] for an incremental design, implementation, and deployment of quantum software services. Systematic mapping studies (SMS) use an evidence-based software engineering approach to systematically identify, analyse, synthesise, and document research trends on a given topic [23]. The academic community leverages QC platforms to advance research and develop prototypes [14], while quantum computing vendors use them to generate revenue and test their evolving quantum systems [15]. By synergising research and development on quantum computing with service-orientation, researchers and developers can apply established design patterns, architectural styles, and modeling techniques to deliver QC as a service [16]. In the context of QSE, this research synthesises the available evidence reflected via published research to provide a software engineering perspective on architecting, implementing, and operationalising QCaaS. We conducted this study to answer three research questions that investigate (i) demographics: trends, types, frequency and progression of research (ii) quantum service lifecycle: that can be integrated in a layered reference architecture for conceptualising, modeling, assembling, and deploying services, and (ii) emerging research: to highlight topics for potentially futuristic research on QCaaS. The results indicate that in quantum service-orientation, QSRs, both functional aspects and quality attributes, guide the four-phase service development lifecycle. Software modeling languages (e.g., UML, circuit diagrams) and patterns (e.g., API gateway, quantum-classic split) can be applied within the proposed layered architecture of QCaaS via quantum-classical hybrid computing. Emerging trends include QSR classification, model-driven development, empirically derived patterns, and defined human roles in QCaaS. The main contributions and implications of this research are to:
@@
This research represents one of the pioneering efforts to rely on evidence-based software engineering guidelines to investigate QCaaS and its implications on QSE. The results can offer academic researchers a mapped overview of existing solutions, open challenges and directions for future QCaaS research [1][8]. Practitioners can also benefit by leveraging insights on patterns (reusability), notations (solution representation), and implementation (prototyping) to design and architect QCaaS solutions [14].
Technical context of QC and QSE is presented in Section 2. Research method is discussed in Section 3. Results of the study to answer the RQs are presented from Section 4 - 7. Review of the existing research is presented in Section 8. Implications and validity threats are discussed in Section 9. Conclusions and future work are presented in Section 10.
@@ 2 Research Context
This section contextualises the core concepts used in our study and elaborates them based on illustrations in Figure 2. First, we overview the fundamental elements of quantum computing and the role of quantum software engineering to operationalise QC systems in Section 2.1, as shown in Figure 2 a). Second, we explain service-orientation from quantum software engineering perspective in Section 2.2 via Figure 2 b). The concepts and terminology introduced in this section are used consistently throughout the paper. They provide a foundational overview of the topic, supported by key references for readers seeking further detail.
Figure 2: An Illustrative Overview of Quantum Computing and Service-Orientation
@@ 2.1.1 Quantum Computing Systems
We explain the QC systems from both hardware and software perspectives, as illustrated in 2 a). This figure, adapted from [19], has been extended to emphasise software engineering and architecture within the context of quantum service-orientation. At the core of quantum computations is the concept of QuBits, the fundamental units of quantum information, which operate by manipulating QuGates [2] [10]. In classical digital systems, binary digits (Bits) are represented as [1, 0], where 1 denotes the ON state and 0 denotes the OFF state, controlling binary gates in digital circuits. In contrast, a QuBit represents a two-state quantum system, expressed as $|0\rangle$ and $|1\rangle$. The state of a single QuBit can be expressed as $|0\rangle=\begin{bmatrix}1\\ 0\end{bmatrix}$ and $|1\rangle=\begin{bmatrix}0\\ 1\end{bmatrix}$ and quantum superposition allows a QuBit to attain a liner combination of both states, expressed as:
@@
As shown in Figure 2 a), a Bit can hold a value of either OFF:0 or ON:1 with 100% probability. In contrast, a QuBit can exist in the state $|0\rangle$, $|1\rangle$, or in a superposition, such as 50% $|0\rangle$ and 50% $|1\rangle$. Additionally, two QuBits can become entangled, meaning that measuring one QuBit instantaneously reveals the state of the other QuBit. For more detailed discussions on QuBits and QuGates in QC systems, please refer to [1] and [14]. To harness quantum computing resources like processors and memory, control software is required to program QuBits and manage QuGates within a QC system. To support systems and services for quantum computing, there is a need for systematic processes that incorporates tools, models, patterns, frameworks, and human expertise as engineering practices to develop quantum software [24] [25].
@@ 2.1.2 Quantum Software Engineering
QSE has emerged as a quantum-specific extension of classical software engineering [26]. QSE integrates the principles of quantum mechanics with software engineering practices to design, develop, validate, and evolve software systems, services, and applications for the quantum era computing [27]. It can enable software designers and developers with structured processes, architectural models for design visualisation, and reusable patterns, enabling them to engineer software that can run effectively on quantum computers. Quantum software systems depend on quantum source code compilers, which allow algorithm designers and programmers to develop and execute software on QCs. For instance, developers can use quantum programming languages like Q# by Microsoft or Qiskit by IBM along with a quantum compiler such as Qiskit transpiler to perform programmable quantum computations [4] [27]. Quantum software systems, capable of managing and controlling quantum hardware, have applications in areas including quantum cryptography, bio-inspired computing, and quantum information processing [1]. However, the limited availability of quantum hardware, the shortage of skilled quantum software professionals, and the high costs of owning or maintaining quantum computers pose significant challenges to the commercial viability of quantum computing [2] [3]. In response, vendors offering QCaaS platforms adopt a service-oriented approach, providing quantum resources to customers as a utility computing model [11] [14].
@@ 2.2.1 Services Computing
It adopts a service-oriented architecture (SOA) approach, enabling users to discover and access a variety of software services that encapsulate computing resources and applications provided by vendors [12]. Fundamental to the concept of service computing is software service that represents a self-contained, reusable, and platform-independent software module that encapsulates specific functionality or business logic, accessible over a network through standardised protocols, typically as part of SOA [28]. More specific types of services can be web services or microservices that enable Software as a Service (SaaS) model in which software applications are hosted and managed by a provider, accessed by users over the network via a web client, and offered on a subscription basis thus eliminating the need for local installation and maintenance [15]. Figure 2 b) illustrates a SOA-based quantum service model, where a QC user (service requester) accesses quantum resources provided by vendors (service providers) through quantum services.